OpenStudy (anonymous):

Bayside insurance offers two health plans. Under plan A,Giselle would have to pay the first $120 of her medical bills, plus 40% of the rest. Under plan B, she would pay the first $140, but only 25% of the rest. For what amount of medical bills would plan B save her money? Assume she has over $140 in bills

OpenStudy (anonymous):

find the break even point by setting the two equations equal. if x is the amount of her medical bills for the first plan she would pay \[120+.40x\] and for the second would pay \[140+.25x\] so write \[120+.40x=140+.25x\] and solve for x via \[.15x=20\] \[x=\frac{12}{.15}=133.3\]

OpenStudy (anonymous):

round to the nearest whole number, and its not 133

OpenStudy (anonymous):

wrong answer

OpenStudy (anonymous):

if you guys do not wanna help,dont screw me over with wrong answers

OpenStudy (anonymous):

Solve the following equation for x:\[130+\frac{40}{100}x=140+\frac{25}{100}x \]x, the break even point, = 200/3 or $66.67 rounded.\[\frac{200}{3}<x \]